OpenStudy (anonymous):

Find the area of the shaded portion intersecting between the two circles. help me i dnt have enough time to get this work done before scool starts so HELP ME!! click pic below

OpenStudy (anonymous):

i cant see a pic

OpenStudy (saifoo.khan):

same here!

OpenStudy (anonymous):

and they gave you no numbers to work with?

OpenStudy (anonymous):

is in the diagram

OpenStudy (anonymous):

This is taking a 60 degree arc area and subtracting the area of an equilateral triangle. We have 2 of these so it's that value times 2. This is 1/6 of a circle area so: \[ \frac{1}{6} \pi (4)^{2} = \frac{8}{3} \pi \] The triangle area is \[\frac{1}{2} base*height\] The height is calculated using pythagoreans. We have a hypotenuse of 4 and a side of 2 so the height is: \[4^{2} - 2^{2} = 16-4 = 12 = 2 \sqrt{3}\] The base of the triangle is 4. The height is 2 sqrt 3. So the area is: \[\frac{1}{2}(4)(2\sqrt{3})=4\sqrt{3}\] The area of the sliver is: \[\frac{8}{3}\pi-2\sqrt{3}\] We have two of these slivers of circle so we will be multiplying the area by 2: \[2(\frac{8}{3}\pi-2\sqrt{3}=\frac{16}{3}\pi-4\sqrt{3}\]